OpenStudy (anonymous):

Find the 10th term of an arithmetic sequence if t1 = 2.1 and t4 = 1.83.

OpenStudy (anonymous):

t1=a=2.1. t4=a+3d=1.83 a+3d=1.83 , 2.1+3d=1.83 , 3d=-0.27 ,d=-0.09 . now u got the (a) and the (d) .. according to the rule :Tn=a+(n-1)d .. T10=2.1+9*-0.09=1.29.

OpenStudy (anonymous):

if u need more explanation tell me ..

OpenStudy (anonymous):

got it! thanksss =)